1 + 2/3 / 1 - 2/3 the symbol in the middle means division the answer is 5 but I dont understand how. Could someone show me step by step.
Thanks in advance!

Answer by Cromlix(4381)

Hi there,
Take the top equation 1 + 2/3
This really 1/1 + 2/3
Common denominator 3
obtained by multiplying
1 and 3
Now cross multiply
3 + 2/3 = 5/3
...............
Take the bottom equation 1 - 2/3
This is really 1/1 - 2/3
Common denominator 3
obtained by multiplying
1 and 3
Now cross multiply
3 - 2/3 = 1/3
................
You now have 5/3 on the top
and 1/3 on the bottom
Because 1/3 is dividing
it is turned upside down
You now have
5/3 x 3/1
3's cancel out
= 5/1 or 5
Hope this helps :-)